OpenStudy (anonymous):

What is the radius of a circle with the equation x^2+y^2+6x-2y+3=0?

OpenStudy (anonymous):

hint: COMPLETE THE DAMN SQUARE! :DDD

OpenStudy (ybarrap):

(x+3)^2+(y-1)^2 -7 = x^2+y^2+6x-2y+3 = 0 So, (x+3)^2+(y-1)^2=7=r^2, where r is the radius Does this make sense?

OpenStudy (anonymous):

no, im confused